Reports 1-1 of 1 Clear search Modify search
AdV-DET (Photodiodes + demodulation boards (longitudinal control))
alloca, bersanetti, gouaty, letendre, masserot, pacaud - 17:28 Monday 09 September 2019 (46872) Print this report
DET - ISC - DAQ: tests changing the VGA gain of the B1, B1p, B2 and B4 RF control PDs

First we changed the VGA gain from high gain mode to low gain mode for the B1, B1p, B2 and B4 longitudinal photodiodes by keeping the same gain value and without re-starting the internal clocks of the mezzanine to keep the demodulation phase.

As today there are 2 types of DaqBox firmware,  we used 2 different ways to update the VGA gain without fully reconfigure the DaqBox

  • for the v4r2 firmware: we used the v14r2 DaqBox server
  • for the v3r0 firmware: we used some expert commands to set the new gain value. As consequence for these DaqBox if someone modifies the demodulation mezzanine configuration, the mezzanine will be reconfigured with  the gain restored  to the high gain mode and the demodulation phase  will have to be adjusted

VGA gain:

  • B2_PD1 and B2_PD2; previous:0x84 - now: 0x1C
  • B4_PD1: previous: 0x89 - now: 0x20 instead of 0x40 to avoid any saturation of the ADC during the lock acquistition
  • B4_PD2: previous: 0x89 - now: 0x40
  • B1p_PD1, B1p_PD2, B1_PD1 and B1_PD2: previous: 0x86 - now 0x2A

With these new gains, the ITF was back at LOW_NOISE_3_SQZ at 2019-09-09 09:11:32UTC-GPS1252055510 (see the attached trend plot). Here the differents sequences of test perfomed

  • From 09h30UTC to 10h05UTC with the ITF in LOW_NOISE_3_SQZ we tried to apply differents 6MHz modulation depth (-28dB to 1dB, 1dB to -28dB) (see this file for the detail)
  • Starting from 10h30UTC  we tried to relock the ITF with the 6MHz modulation always present  and by  adjusting the 6MHz depth at ITF step 120 from 1dB to -1dB.  With -1dB  the ITF succeeded to lock at  LOW_NOISE_3_SQZ
  • The last trial, from 13h04UTC to 13h20UTC  the ITF was relocked at LOW_NOISE_3_SQZ with a 6MHz modulation depth of -1dB , at 13h20UTC the 6MHz modulation depth was set to  -28dB . The TF unlocks at 13h40UTC

 

 

Images attached to this report
Non-image files attached to this report
Comments to this report:
tacca - 18:28 Tuesday 10 September 2019 (46883) Print this report
Since the change of the VGA gains, the values of the B4_112MHz signals have decreased a lot.
Images attached to this comment
mantovani - 11:42 Wednesday 11 September 2019 (46889) Print this report

This morning, while I was trying to optimize the AA working point I could not increase this value. Can be checked if there is a gain in the PDs that is not as before?

 

masserot, gouaty - 15:40 Wednesday 11 September 2019 (46896) Print this report

No update were performed on the quadrant demodulated channels

masserot, gouaty - 15:48 Wednesday 11 September 2019 (46893) Print this report

Thank to the observations made by  Matteo, we have made more measurements to characterize the VGA gain update from high gain to low gain in the case of the B4 signals

The setup  is the following

  • Demod board CH00 input:  0.1V@112MHz  demodulated at 112MHz
  • Demod board CH02 input : 0.1V@56MHz  passing through a selective filter @56MHz demodulated at 56MHz
  • An Acl server to compute the magnitude of the demodulated signals

The attached plot show  for each input the FFT of the channel acquired at 400MHz, the ADC maximun value and minimal value multiplied by -1  and the magnitude of the extracted line.

One clearly see that the gain modification between 0x89 HG to 0x40 LG is larger than the expected difference of 0.45%

The following table show the measured magnitude for the 2 modes of gain

As consequence all the B4 demodulated channels are lower than previously by 7% for the 112MHz, by  4.6% for the 56MHz

 The correction factors can be applied easily only for the B4 signals by

  •  in  the SPRB_dbox_bench configuration file , updating for each  DBOX_DEMOD_SIGNAL the slope parameter then reloading the configuration
  • and then reloading the SSFS_Ctrl configuration file
Images attached to this comment
tacca - 15:50 Wednesday 11 September 2019 (46898) Print this report
In the attached figure, the images acquired by the phase cameras when the 6 MHz modulation depth was set at 1dB.
Images attached to this comment
Search Help
×

Warning

×